# Bifurcated Attention for Single-Context Large-Batch Sampling

### v1

@@ Abstract
In our study, we present bifurcated attention, a method developed for language model inference in single-context batch sampling contexts. This approach aims to reduce redundant memory IO costs, a significant factor in latency for high batch sizes and long context lengths. Bifurcated attention achieves this by dividing the attention mechanism during incremental decoding into two distinct GEMM operations, focusing on the KV cache from prefill and the decoding process. This method ensures precise computation and maintains the usual computational load (FLOPs) of standard attention mechanisms, but with reduced memory IO. Bifurcated attention is also compatible with multi-query attention mechanism known for reduced memory IO for KV cache, further enabling higher batch size and context length. The resulting efficiency leads to lower latency, improving suitability for real-time applications, e.g., enabling massively-parallel answer generation without substantially increasing latency, enhancing performance when integrated with post-processing techniques such as reranking.
@@ 1 Introduction
The advent of large language models (LLMs) has ushered in a new era of machine learning, exhibiting remarkable performance on a wide array of tasks (Brown et al., 2020; OpenAI, 2023; Chowdhery et al., 2022; Touvron et al., 2023; Chen et al., 2021; Hoffmann et al., 2022; Li et al., 2022; Microsoft, ; Amazon, 2022; Nijkamp et al., 2023). Despite their impressive capabilities, the deployment of these large-scale models in practical applications poses significant challenges, particularly in terms of inference latency and efficiency. Enhancing these aspects is critical, as they directly influence the computational resources required to generate predictions and enable the practical implementation of these advanced models across various industries.
@@ B.1 Applications of Single-Context Batch Sampling
  • Chatbots and Conversational AI: Conversational agents often need to generate multiple responses to handle different interpretations of a user’s input or to provide multiple suggestions. The reduced latency offered by the proposed method can significantly improve the responsiveness of chatbots, leading to a more natural and fluid conversation with users (Google, 2023).
@@ B.2 Supporting Long Context Requires IO-Efficient Attention
As language models are becoming general purpose and highly capable, the demand for language models to handle longer context sequences has grown significantly. Recently, there is an ongoing focus on models that can handle even longer context sequences (Bulatov et al., 2023; OpenAI, 2023; Team, 2023). As of today, GPT-4 (OpenAI, 2023) supports context length of 32k tokens, and MPT-7B (Team, 2023) extends it to 64k while Anthropic’s Claude  [^3] https://www.anthropic.com/index/100k-context-windows supports as long as 100k input length. Most recently, Bulatov et al proposed 1M token input context length for transformers. These models push the boundaries of context understanding and generation capabilities, enabling more comprehensive discourse understanding and contextually informed responses.
@@ C.1 Model Training Details
We use AdamW optimizer (Kingma and Ba, 2014) with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=10^{-8}$. The warm-up steps were set to 2000, and a cosine annealing learning rate schedule was employed after reaching the peak learning rate. The minimum learning rate was set to 10% of the corresponding peak learning rate. A weight decay (Loshchilov and Hutter, 2017) of 0.01 and gradient clipping of 1.0 were applied to enhance training stability. Following the approach in (Shoeybi et al., 2019), the standard deviation for random weight initialization was rescaled for larger models. Our training pipeline is based on PyTorch Lightning and we use bfloat16 (Kalamkar et al., 2019) and DeepSpeed (Rasley et al., 2020) for training optimization. Finally, a random split of 0.1% of the data was reserved as a validation set.
@@ C.2 Model Configurations
Model Family
Attention Type
$groups$
$d_{head}$
$n_{layer}$
$N_{params}$ (billions)
125M
MH
12
64
12
0.125
MG
4
64
12
0.115
MQ
1
64
12
0.112
672M
MH
20
72
24
0.672
MG
4
72
24
0.592
MG + $2\times d$
4
72
24
0.393
MQ
1
72
24
0.578
2.8B
MH
24
128
24
2.878
MG
4
128
24
2.501
MG + $2\times d$
4
128
24
1.595
MQ
1
128
24
2.444
13B
MH
40
128
40
12.852
MG
8
128
40
11.174
MQ
1
128
40
10.807
@@ C.3 Model Details of 1B Latency Experiment
Model Family
Attention Type
$groups$
$d_{head}$
$n_{layer}$
$N_{params}$ (billions)
1B
MH
20
128
12
1.077
MG
4
128
15
1.156
MQ
1
128
16
1.193

### v2

# Bifurcated Attention: Accelerating Massively Parallel Decoding with Shared Prefixes in LLMs

@@ Abstract
This study introduces bifurcated attention, a method designed to enhance language model inference in shared-context batch decoding scenarios. Our approach addresses the challenge of redundant memory IO costs, a critical factor contributing to latency in high batch sizes and extended context lengths. Bifurcated attention achieves this by strategically dividing the attention mechanism during incremental decoding into two separate GEMM operations: one focusing on the KV cache from prefill, and another on the decoding process itself. While maintaining the computational load (FLOPs) of standard attention mechanisms, bifurcated attention ensures precise computation with significantly reduced memory IO. Our empirical results show over 2.1$\times$ speedup when sampling 16 output sequences and more than 6.2$\times$ speedup when sampling 32 sequences at context lengths exceeding 8k tokens on a 7B model that uses multi-head attention. The efficiency gains from bifurcated attention translate into lower latency, making it particularly suitable for real-time applications. For instance, it enables massively parallel answer generation without substantially increasing latency, thus enhancing performance when integrated with post-processing techniques such as re-ranking.
@@ 1 Introduction
The advent of large language models (LLMs) has ushered in a new era of machine learning, exhibiting remarkable performance on a wide array of tasks (Brown et al., 2020; OpenAI, 2023; Chowdhery et al., 2022; Touvron et al., 2023; Chen et al., 2021; Hoffmann et al., 2022; Li et al., 2022; Microsoft, ; Amazon, ; Nijkamp et al., 2023). Despite their impressive capabilities, the deployment of these large-scale models in practical applications poses significant challenges, particularly in terms of inference latency and efficiency. Enhancing these aspects is critical, as they directly influence the computational resources required to generate predictions and enable the practical implementation of these advanced models across various industries.
@@ 2 Related Work
Additionally, we acknowledge concurrent work by Juravsky et al. (2024) which presents methods to improve inference efficiency with shared-prefixes, that coincides with bifurcated attention.
@@ 5.2.2 Single-Context Batch Sampling
5.3 Compatibility with Torch-Compile
Bifurcated attention can be implemented with 4 einsum calls in native PyTorch, making it compatible with Torch-Compile. With Torch Compile, we can take advantage of kernel-fusion and concurrency to improve the latency of the model. To demonstrate this, we implement bifurcated attention on top of GPTFast (PyTorch, 2023)  [^3] Link to our code: https://github.com/bifurcated-attn-icml-2024/gpt-fast-parallel-sampling.
We experiment on a 7B parameter model, with a hidden dimension of 4096 that is 32 layers deep and has 32 heads, as shown in Table LABEL:tab:torch-compile. We observe that the overall latency imporvements fwith large-batch sampling with respect to standard SDPA attention remains consistent even when we compile the model.
Table 1: Per-token latency (ms) of a 7B multi-head model on GPT-Fast with and without Torch Compilation compared to a model using Torch’s standard SDPA kernel.
without Compile Compiled
Context BS SDPA Bifurcated SDPA Bifurcated
8k 1 26.40 30.39 8.78 8.64
2 28.71 31.37 10.51 11.77
4 43.36 31.44 13.23 12.03
8 72.71 33.72 17.33 12.36
16 132.89 31.71 26.19 12.60
16k 1 30.13 30.66 13.06 12.16
2 44.74 32.62 15.35 17.17
4 73.62 33.44 20.65 17.33
8 132.29 34.67 32.06 18.07
16 251.47 36.78 OOM 18.46
32k 1 44.94 39.97 19.80 20.90
2 69.22 48.61 OOM 29.34
4 OOM 49.77 - 29.73
8 - 51.31 - 30.30
16 - 54.92 - 30.66
@@ 5.4 Applications
Many reasoning algorithms such as self-consistency Chain-of-thought (SC-COT) (Wang et al., 2023) and Tree-of-thought (ToT) (Yao et al., 2023) depend on sampling multiple outputs with shared prefix, where bifurcated attention will enable higher accuracy under same costs.
@@ 6 Conclusion
Impact Statement
Bifurcated attention is an approach that can significantly reduce the latency and associated costs involved in deploying large language models (LLMs). A key advantage of this technique is its potential to lower the carbon emissions associated with LLM inference. While reducing deployment costs could potentially lead to broader adoption of LLMs, the societal impact of such increased usage remains difficult to predict with certainty. Nonetheless, bifurcated attention presents an opportunity to make LLM deployment more efficient and environmentally friendly, although the broader implications warrant careful consideration.
@@ B.1 Applications of Single-Context Batch Sampling
  • Chatbots and Conversational AI: Conversational agents often need to generate multiple responses to handle different interpretations of a user’s input or to provide multiple suggestions. The reduced latency offered by the proposed method can significantly improve the responsiveness of chatbots, leading to a more natural and fluid conversation with users (Google, ).
@@
  • Reasoning: using Self-consistency Chain-of-Thought (CoT-SC) (Wang et al., 2023) and Tree-of-Thought (ToT) (Yao et al., 2023) requires the model to sample multiple outputs with a shared prefix. Bifurcated attention will enable larger number of reasoning paths in SC-COT and larger trees in ToT at the same cost of inference.
@@ B.2 Supporting Long Context Requires IO-Efficient Attention
As language models are becoming general purpose and highly capable, the demand for language models to handle longer context sequences has grown significantly. Recently, there is an ongoing focus on models that can handle even longer context sequences (Bulatov et al., 2023; OpenAI, 2023; Team, 2023, 2023). As of today, GPT-4 (OpenAI, 2023) supports context length of 32k tokens, and MPT-7B (Team, 2023) extends it to 64k while Anthropic’s Claude  [^4] https://www.anthropic.com/index/100k-context-windows supports as long as 100k input length. Most recently, Bulatov et al proposed 1M token input context length for transformers. These models push the boundaries of context understanding and generation capabilities, enabling more comprehensive discourse understanding and contextually informed responses.
@@ C.1 Model Training Details
We use AdamW optimizer ((Kingma and Ba, 2014)) with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=10^{-8}$. The warm-up steps were set to 2000, and a cosine annealing learning rate schedule was employed after reaching the peak learning rate. The minimum learning rate was set to 10% of the corresponding peak learning rate. A weight decay ((Loshchilov and Hutter, 2017)) of 0.01 and gradient clipping of 1.0 were applied to enhance training stability. Following the approach in ((Shoeybi et al., 2019)), the standard deviation for random weight initialization was rescaled for larger models. Our training pipeline is based on PyTorch Lightning and we use bfloat16 ((Kalamkar et al., 2019)) and DeepSpeed ((Rasley et al., 2020)) for training optimization. Finally, a random split of 0.1% of the data was reserved as a validation set.
@@ C.2 Model Configurations
Model Family Attention Type $groups$ $d_{head}$ $n_{layer}$ $N_{params}$ (billions)
125M MH 12 64 12 0.125
MG 4 64 12 0.115
MQ 1 64 12 0.112
672M MH 20 72 24 0.672
MG 4 72 24 0.592
MG + $2\times d$ 4 72 24 0.393
MQ 1 72 24 0.578
2.8B MH 24 128 24 2.878
MG 4 128 24 2.501
MG + $2\times d$ 4 128 24 1.595
MQ 1 128 24 2.444
13B MH 40 128 40 12.852
MG 8 128 40 11.174
MQ 1 128 40 10.807
@@ C.3 Model Details of 1B Latency Experiment
Model Family Attention Type $groups$ $d_{head}$ $n_{layer}$ $N_{params}$ (billions)
1B MH 20 128 12 1.077
MG 4 128 15 1.156
MQ 1 128 16 1.193
@@ Appendix G Compatibility with Speculative Decoding and Fast Decoding techniques
Table 6: Per-token generation latency (ms) with bifurcated attention compared to native Flash attention 2 kernel and Torch’s SDPA attention kernel implementations, with and without using the torch compile option. Measurements are taken using a 7B parameter model (32 layers, 32 heads, hidden dimension = 4096) with multi-head attention. SDPA Math represents the default attention operations by Torch, while SDPA Flash utilizes Flash attention under the hood. "NC" refers to the use of non-contiguous memory allocation for the cache, allowing reuse of the cache from the prompt. Note that Flash attention kernels are currently not compatible with torch-compile. The experiment results below utilize an Nvidia H100 GPU.
without Torch Compile with Torch Compile
BS Bifurcated Flash2 SDPA Math SDPA Flash Flash2 SDPA Flash SDPA Math Bifurcated SDPA Math
(NC) (NC) (NC)
Context Length : 8k
1 30.38 24.06 26.39 22.00 24.54 23.43 10.66 8.63 8.77
2 31.37 24.49 28.70 24.77 31.53 31.66 14.45 11.74 10.50
4 31.44 39.66 43.36 38.86 50.54 51.06 23.20 12.03 13.22
8 33.72 60.92 72.70 61.22 84.52 84.99 35.42 12.36 17.33
16 31.70 109.64 132.89 109.45 155.85 159.82 63.68 12.59 26.19
32 31.78 205.57 251.02 205.92 305.39 306.60 120.39 13.47 -
64 35.26 OOM OOM - 599.08 601.48 238.19 15.35 -
128 48.69 - - - 1183.46 OOM OOM 19.56 -
256 75.21 - - - 1842.98 - - 27.15 -
512 130.58 - - - - - - 44.33 -
1024 242.73 - - - - - - 81.14 -
2048 473.74 - - - - - - OOM -
Context Length : 16k
1 30.66 26.28 30.13 26.22 30.49 30.20 15.53 12.16 13.06
2 32.62 37.72 44.74 38.25 51.30 51.24 22.46 17.17 15.35
4 33.44 65.98 73.62 65.83 91.25 90.76 39.51 17.33 20.65
8 34.67 110.31 132.29 110.55 159.96 160.39 64.22 18.07 32.06
16 36.78 206.93 251.47 206.52 306.75 307.31 119.87 18.46 OOM
32 41.93 OOM OOM OOM 601.10 603.61 237.89 19.92 -
64 50.53 - - - 1195.35 OOM OOM 22.96 -
128 68.31 - - - 1908.23 - - 28.98 -
256 106.10 - - - OOM - - 40.07 -
512 183.14 - - - - - - 65.02 -
1024 339.74 - - - - - - 117.75 -
2048 660.20 - - - - - - OOM -
Context Length : 32k
1 39.97 37.67 44.94 37.46 67.44 67.30 30.39 20.90 19.80
2 48.61 55.94 69.22 55.86 156.61 156.35 47.63 29.34 OOM
4 49.77 OOM OOM OOM 300.47 300.97 90.19 29.73 -
8 51.31 - - - 567.93 568.81 152.19 30.30 -
16 54.92 - - - 670.21 672.42 290.59 30.66 -
32 62.28 - - - 1318.05 1323.25 569.74 32.15 -
64 75.22 - - - OOM OOM OOM 35.25 -
128 101.18 - - - - - - 41.44 -
256 159.09 - - - - - - OOM -
512 277.05 - - - - - - - -
1024 OOM - - - - - - - -
Appendix H Experiments with GPTFast
The implementation of context-aware bifurcated attention in native PyTorch demonstrates significant reductions in parallel sampling latency for both multi-headed attention (MHA) and grouped query attention (GQA) architectures. Bifurcated attention, being context-aware and implemented natively in PyTorch, can directly benefit from PyTorch’s compilation capabilities.
We observe Bifurcated attention outperforming FlashAttention2, especially for larger context lengths and higher degrees of tensor parallelism. Since Bifurcated attention is primarily targeting decode phase during inference, leveraging the efficiency of FlashAttention2 for the prefill (context encoding) step.
H.1 In Comparison with FlashAttention
Table 7: Per-token generation latency (ms) with bifurcated attention compared to the native Flash attention kernel. Measurements are taken with a 7B parameter model (32 layers, 32 heads, hidden dimension = 4096, 8 kv heads) using grouped query attention. Note that Flash attention kernels are currently not compatible with torch-compile. In this table, "NC" refers to the use of non-contiguous memory allocation for the cache, allowing reuse of the cache from the prompt. The experiment results below utilize an Nvidia H100 GPU.
BS Bifurcated + Compile Bifurcated Flash2 Flash 2 (NC)
Context: 8k 16k 32k 8k 16k 32k 8k 16k 32k 8k 16k 32k
1 10.56 15.16 22.79 28.37 30.97 37.20 21.76 23.59 26.64 23.48 25.23 28.20
2 11.35 15.99 23.72 29.53 32.16 37.47 22.46 23.78 26.82 39.93 28.53 45.70
4 11.52 16.20 23.98 29.58 32.19 37.48 22.57 24.22 27.30 71.57 42.47 72.94
8 11.79 16.61 24.59 29.58 32.41 38.12 22.65 24.03 28.36 126.35 70.01 127.96
16 11.72 16.68 24.87 30.27 32.85 37.29 22.31 30.19 OOM 240.96 130.77 245.81
32 12.50 17.77 27.01 29.76 32.75 37.84 26.06 OOM 468.93 244.54 467.61
64 13.87 19.90 30.31 29.52 32.07 45.73 OOM 403.08 482.71 463.55
128 17.03 24.90 37.60 29.55 40.26 63.06 788.66 465.70 909.02
256 24.38 33.76 52.06 40.07 59.42 96.28 915.89 1805.60
512 39.08 OOM OOM 65.74 OOM OOM OOM OOM
1024 72.24 118.57
2048 OOM 230.88
FlashAttention is a highly efficient general-purpose fused attention kernel that is particularly effective during context encoding, as it avoids materializing the expensive-to-read-and-write $n\times n$ attention matrix in GPU memory.
However, during incremental decoding with single-context batch sampling, native FlashAttention kernels are not as efficient because they are not designed to be context-aware. Specifically, if there are B batch indices of K,V cache that are duplicate in values due to the shared prefix, FlashAttention-2 (FA2) can use paged KV-Cache to refer and point them to the same KV-pairs for the prefix across a batch. Nevertheless, this does not prevent the FlashAttention kernel from performing multiple reads of the KV-pairs from the shared prefix.
Table 6 shows that a context-aware approach such as bifurcated attention outperforms FlashAttention in parallel sampling scenarios, especially with an increasing number of parallel samples. Notably, the bifurcated attention kernel is utilized solely during the decode step, allowing the efficient FlashAttention2 kernel to be employed during the prefill step for context lengths up to 8192. Furthermore, while non-contiguous memory avoids out-of-memory issues during parallel sampling for non-context-aware kernels, bifurcated attention’s memory setup, which maintains only one copy of the context and expands by reference across batch indices, achieves substantially lower latencies. However, the native FlashAttention2 implementation is not yet compatible with PyTorch’s compilation capabilities.
In the future, it may be possible to combine bifurcated attention with FlashAttention to optimize the latency further.
H.2 Trends with Grouped Querry Attention (GQA)
For GQA architectures, bifurcated attention is able to help scale to very large inference workloads. Using PyTorch’s compilation mode, the inference with bifurcated attention is much faster compared to FlashAttention2. Table 7 presents the results for context lengths of 8K, 16K, and 32K tokens. Note that PyTorch’s SDPA is not directly supported for GQA and thus not included in the comparison.
H.3 Compatibility with Tensor Parallel (TP)
Higher tensor parallelism is often required to handle higher inference workloads, as seen in Table 8. The proposed context-aware bifurcated attention method works out-of-the-box without additional modifications for tensor parallelism. With TP we get to work with much larger context lengths.
Table 8: Per-token generation latency (ms) Mistral 7B at different context lengths with TP=2 experimented on Nvidia’s H100 GPU.
Context BS SDPA Bifurcated Flash2
16384 16 131.46 55.51 92.11
32640 8 133.85 58.56 92.35
32640 16 246.53 58.00 162.02
32640 32 OOM 57.86 OOM
32640 64 60.33
32640 128 67.82